Please help me solve this problem: the difference between the measures of two complementary angles is 14. Find both angles.
Answer by nerdybill(6963)

Please help me solve this problem: the difference between the measures of two complementary angles is 14. Find both angles.
.
Two angles are complementary if the sum of the two angles equals 90 degrees.
.
Let x = measure of one angle
then
90-x = measure of second angle
.
x - (90-x) = 14
x - 90+x = 14
2x - 90 = 14
2x = 104
x = 52 degrees
.
second angle:
90-x = 90-54 = 36 degrees
.
answer: 52 and 36 degrees